Divide. Write result in simplest form:
3/4 divide by 9

5 1/3 divided by 2 2/5
2 problems to resolve
Thanks

Answer by fractalier(6550)

3/4 divided by 9 is equal to 3/4 times 1/9 = 3/36 = 1/12
and
5 1/3 divided by 2 2/5 = 16/3 divided by 12/5 = 16/3 * 5/12 = 80/36 = 20/9 = 2 2/9